# Flexible MOF Generation with Torsion-Aware Flow Matching

### v1

@@ 6 Experiments
We evaluate MOFFlow-2 on two key tasks, MOF structure prediction and MOF generation, to demonstrate its ability to predict accurate 3D structures and design novel MOFs. We first describe the shared data preprocessing pipeline, then present the structure prediction task in Section 6.1 and the MOF generation task in Section 6.2.
Data preprocessing. We generally follow the preprocessing pipeline from prior work [11, 10]. Starting with the dataset from Boyd et al. [36], we apply metal-oxo decomposition algorithm from MOFid [37] and discard any structures containing more than 20 building blocks [11]. The resulting dataset is split into an 8:1:1 ratio for train/valid/test sets in the structure prediction task, and into a 9.5:0.5 train/valid split for MOF generation [11, 10]. Since the dataset [36] consists of hypothetical structures, we further filter out invalid MOFs using MOFChecker [38], and apply the MOF matching procedure to address distributional shift at test time, as discussed in Section 5.3. Preprocessing details and dataset statistics are available in Appendix F.
@@ 6.1 MOF structure prediction
Baselines. We compare our model against both classic optimization- and learning-based approaches. For the classic baselines, we consider random search (RS) and the evolutionary algorithm (EA) implemented in CrySPY [39]. Among the learning-based baselines, we include DiffCSP [8], a general crystal structure prediction model, and MOFFlow [10], a MOF-specific structure prediction model. To ensure a fair comparison, we assume that no models have access to ground-truth local coordinates. Accordingly, we retrain MOFFlow with matched coordinates (Section 5.3) and evaluate it using the initialized structures described in Section 4.2.
@@
Results. Table 2 shows that MOFFlow-2 outperforms all baselines across both tolerance settings. Consistent with prior work [10], we observe that the optimization-based methods (RS and EA) and the general CSP baseline (DiffCSP) achieve near-zero match rates, indicating their limited effectiveness on MOF structures. Compared to MOFFlow [10], our method achieves higher match rates and lower RMSE, underscoring the importance of explicitly modeling torsion angles for accurate structure prediction. We also provide visualizations of property distributions in Appendix G.
Table 2: Structure prediction accuracy. We compare the structure prediction performance of random search (RS), evolutionary algorithm (EA), DiffCSP, MOFFlow, and MOFFlow-2. MR is the match rate, RMSE is the root mean squared error, and -- indicates no match. $\operatorname{stol}$ is the site tolerance for matching criteria.
@@ 6.2 MOF generation
We evaluate the generative performance of MOFFlow-2 by (1) measuring validity, novelty, and uniqueness (VNU), and (2) comparing the distribution of MOF properties against that of the training set. We benchmark against MOFDiff [11], a coarse-grained diffusion model for MOF generation. Both models generate 10,000 samples using their respective pipelines. To isolate the effect of the generative model, no force-field relaxation is applied to the generated structures.
@@
Results. Table 3 summarizes the generation performance of MOFFlow-2 compared to MOFDiff. MOFFlow-2 achieves higher scores across all metrics. Notably, MOFFlow-2 not only generates MOFs with novel combinations of known building blocks but also produces entirely new building blocks, demonstrating its potential for discovering previously unseen MOFs. In addition, MOFFlow-2 is also faster than MOFDiff, which depends on an optimization-based self-assembly procedure for final structure construction. As shown in Figure 4, the property distribution of MOFFlow-2 aligns more closely with the ground-truth. Importantly, it spans a broader range of values than MOFDiff, indicating that MOFFlow-2 generates MOFs with more diverse physical properties. We provide visualizations of representative VNU samples generated by MOFFlow-2 in Figure 5.
@@ 7 Conclusion
We introduced MOFFlow-2, a two-stage generative model for MOF design and structure prediction. In the first stage, a building block generator produces MOF sequences in SMILES format, which are initialized into 3D structures using a predefined metal library and cheminformatics tools. In the second stage, a structure prediction module assembles the complete MOF by jointly predicting rotations, translations, torsion angles, and lattice parameters. Experimental results show that MOFFlow-2 outperforms existing models in both generative design and structure prediction. Our limitation is that we did not thoroughly verify the stability of the generated structures with DFT calculation.
@@ Appendix A Implementation details for building block generator
We provide additional details for implementing the building block generator (Section 4.1), including the tokenization process, batching strategy, and hyperparameter settings.
@@
We illustrate the tokenization process for a canonical MOF sequence defined in Section 4.1, which transforms the string into a sequence of discrete symbols $[b_{1},\dots,b_{S}]$.
@@
We then map each token to a unique vocabulary index to get $[b_{1},\dots,b_{S}]$.
@@
Hyperparameters. In Table 4, we present the hyperparameters for training the building block generator, including the batch configuration, model architecture, and optimizer settings.
@@ Appendix B Analysis and visualization of metal building blocks
Figure 6(a) and Figure 6(b) show the RMSD histograms of the extracted metal building blocks for the generation and structure prediction tasks, respectively. There are only 7 and 8 distinct metal building block types, despite the large sizes of the training datasets (>150k). The RMSD histograms further indicate that all metal types exhibit low structural variability. These observations justify our template-based strategy for initializing the metal structures, described in Section 4.2. We also provide visualizations of the metal building blocks in Figure 7.
@@ C.1 Conditional flows and conditional vector fields
where $\exp:\mathfrak{so}(3)\rightarrow{}SO(3)$ and $\log:SO(3)\rightarrow{}\mathfrak{so}(3)$ are the exponential and logarithmic maps for $SO(3)$, respectively [44]. Since we target $\Delta q_{t\rightarrow{}1}:=q_{t}q_{1}^{\top}$ (Section 5.1),
@@ C.3 Training and inference algorithm
Algorithm 2 and Algorithm 3 present the training and inference algorithms for the structure prediction model, respectively. During training, the model outputs rotations $\Delta\hat{\bm{Q}}_{t\rightarrow 1}\in\mathbb{R}^{M\times 3\times 3}$ and torsions $\hat{\bm{\Phi}}_{1}\in\mathbb{R}^{P\times 2}$ in the Euclidean space, which are directly supervised with the targets $\Delta\bm{q}_{t\rightarrow 1}$ and $\bm{\phi}_{1}$. At inference, these outputs are projected onto their respective manifolds as $\Delta\hat{\bm{q}}_{t\rightarrow 1}=\operatorname{Procrustes}(\Delta\hat{\bm{Q}}_{t\rightarrow 1})\in SO(3)^{M}$ and $\hat{\bm{\phi}}_{1}=\hat{\bm{\Phi}}_{1}/\|\hat{\bm{\Phi}}_{1}\|_{2}\in\mathbb{T}^{P}$.
@@
3:     Sample $\Delta\bm{q}_{1\rightarrow{}0}$, $\bm{\tau}_{0}$, $\bm{\phi}_{0}$, $\bm{\ell}_{0}$ from prior distributions defined in Section 5.1.
4:     Interpolate $\Delta\bm{q}_{1\rightarrow{}t},\bm{\tau}_{t},\bm{\phi}_{t},\bm{\ell}_{t}$ according to Section C.1
5:     Apply $\bm{X}_{t}\leftarrow{}(\Delta\bm{q}_{1\rightarrow{}t},\bm{\tau}_{t},\bm{\phi}_{t})\cdot\bm{X}_{1}$ according to Section C.2.
@@
7:     Optimize loss $\mathcal{L}(\theta)$ in Equation 5.
@@
1:Initialize structure $\tilde{\bm{X}}_{0}\leftarrow{}\operatorname{Initialize}(\mathcal{B}_{\text{2D}})$ according to Section 4.2.
2:Sample $\Delta\bm{q}_{0}$, $\bm{\tau}_{0}$, $\bm{\phi}_{0}$, $\bm{\ell}_{0}$ from prior distributions defined in Section 5.1.
@@ D.1 Initialization module
Building block index. We define the building block index $k$ as introduced in Section 5.2 to resolve ambiguity between building blocks of the same type – i.e., those sharing the same SMILES representation. To assign consistent indices, we apply the following lexicographic ordering rules:
@@ D.2 Interaction module
The interaction module follows a Transformer encoder architecture [46] (Figure 2). Here, we provide further implementation details, including the model architecture (Algorithm 4) and associated hyperparameters (Table 5).
@@
6:$H\leftarrow{}H+\operatorname{MHA}(\operatorname{RMSNorm}(H),E,\mathcal{E})$triangleright Refer to Section 5.2 for details on $\operatorname{MHA}$.
7:$H^{\prime}\leftarrow{}H+\operatorname{FFN}(\operatorname{RMSNorm}(H))$
@@ D.3 Output module
We describe the architectures of the four prediction heads in the output module (Section 5.2). We begin with the block attention pooling module, which aggregates atom-level node embeddings from the interaction module into block-level embeddings (Algorithm 5). We then detail the architectures of the rotation (Algorithm 6), translation (Algorithm 7), lattice (Algorithm 8), and torsion heads (Algorithm 9). The corresponding hyperparameters are summarized in Table 6.
@@
1:# Construct building block coordinates $[x_{\text{bb}}^{(m)}]_{m=1}^{M}\in\mathbb{R}^{M\times 3}$ and features $[h_{\text{bb}}^{(m)}]_{m=1}^{M}\in\mathbb{R}^{M\times D}$.
@@
8:Construct edge distance features $[e_{mj}]_{(m,j)\in\mathcal{E}_{\text{bb}}}$ where $e_{mj}=\operatorname{RBF}({\|x_{\text{bb}}^{(m)}-x_{j}\|_{2}})$.
@@
2:Average and MLP ~$\tilde{\bm{\ell}}_{1}\leftarrow{}\operatorname{MLP}(\sum_{m=1}^{M}h_{\text{bb}}^{(m)})$. $\triangleright$ Two-layer MLP with GELU activation.
@@
1:# Construct rotatable bond coordinates $[x_{\text{rot}}^{(p)}]_{p=1}^{P}\in\mathbb{R}^{P\times 3}$ and features $[h_{\text{rot}}^{(p)}]_{p=1}^{P}\in\mathbb{R}^{P\times D}$.
@@
9:Construct edge distance features $[e_{pj}]_{(p,j)\in\mathcal{E}_{\text{rot}}}$ where $e_{pj}=\operatorname{RBF}(\|x_{\text{rot}}^{(p)}-x_{j}\|_{2})$.
@@
18:Apply $\bm{\Phi}_{1}\leftarrow{}\operatorname{MLP}(H_{\text{rot}})$ where $H_{\text{rot}}=[h_{\text{rot}}^{(p)}]_{p=1}^{P}.$triangleright Two-layer MLP with GELU activation.
@@ Appendix E MOF matching
We describe the MOF matching procedure, a preprocessing step used to reduce distributional shift between training and inference, as outlined in Section 5.3. Given an MOF structure and its building blocks, we first determine whether each block contains metal elements to classify it as metal or organic. For metal blocks, we retrieve the corresponding template from the metal library and align it to the ground-truth structure. For organic blocks, we initialize the structure using RDKit, optimize torsion angles via differential evolution to minimize RMSD with the ground-truth [16], and then align the result. After processing all blocks, we compute the RMSD between the reconstructed and ground-truth structures. This procedure is repeated three times, and structures with final RMSD below 0.5Å are retained. The full algorithm for a single matching iteration is shown in Algorithm 10.
@@
Output: Matched MOF structure $\tilde{\mathcal{S}}=(\bm{A},\tilde{\bm{X}},\bm{\ell})$, $\operatorname{RMSD}(\mathcal{S},\tilde{\mathcal{S}})\in\mathbb{R}_{+}$.
@@
4:         Retrieve the corresponding template structure $\tilde{X}^{(m)}$ from metal library (Section 4.2).
@@
8:         Initialize structure $\tilde{X}^{(m)}$ with RDKit (Section 4.2).
@@ Appendix F Experimental details
Data preprocessing. We detail the data preprocessing steps described in Section 6. First, we discard MOF structures with more than 20 building blocks since large structures that are too large may be difficult to synthesize [11]. Next, we extract key features from each structure, including Cartesian coordinates, RDKit-derived atomic features, Niggli-reduced cells, symmetrically equivalent coordinates for each building block (Section 5.3), and canonical atoms defining the torsion angles (Section 5.3). We then construct the metal library (Section 4.2) and perform MOF matching (Appendix E) on the training dataset. Finally, since the dataset is synthetic [36], we filter out invalid MOFs using MOFChecker [38], which ensures the presence of key elements (e.g., C, H, and a metal), no atomic overlaps or coordination issues, sufficient porosity, and the absence of highly charged fragments or isolated molecules.
Data statistics. We present the data statistics for structure prediction (Section 6.1) in Tables 7, 8 and 9 and those for generation (Section 6.2) in Tables 10 and 11.
Property (number of samples $=157,474$) Min Mean Max
@@
Property (number of samples $=19,603$) Min Mean Max
@@
Property (number of samples $=19,792$) Min Mean Max
@@
Property (number of samples $=187,047$) Min Mean Max
@@
Property (number of samples $=10,243$) Min Mean Max

### v3

@@ 6 Experiments
We evaluate MOFFlow-2 on two key tasks, MOF structure prediction and MOF generation, to demonstrate its ability to predict accurate 3D structures and design novel MOFs. We first describe the shared data preprocessing pipeline, then present the structure prediction task in Section˜6.1 and the MOF generation task in Section˜6.2. Additional experimental details are provided in Appendix˜F.
Data preprocessing. We generally follow the preprocessing pipeline from prior work [11, 10]. Starting with the dataset from Boyd et al. [36], we apply metal-oxo decomposition algorithm from MOFid [37] and discard any structures containing more than 20 building blocks [11]. The resulting dataset is split into an 8:1:1 ratio for train/valid/test sets in the structure prediction task, and into a 9.5:0.5 train/valid split for MOF generation [11, 10]. Since the dataset [36] consists of hypothetical structures, we further filter out invalid MOFs using MOFChecker [38], and apply the MOF matching procedure to address distributional shift at test time, as discussed in Section˜5.3. Preprocessing details and dataset statistics are available in Appendix˜F.
@@ 6.1 MOF structure prediction
Baselines. We compare our model against both classic optimization- and learning-based approaches. For the classic baselines, we consider random search (RS) and the evolutionary algorithm (EA) implemented in CrySPY [39]. Among the learning-based baselines, we include DiffCSP [8], a general crystal structure prediction model, and MOFFlow [10], a MOF-specific structure prediction model. To ensure a fair comparison, we assume that no models have access to ground-truth local coordinates. Accordingly, we retrain MOFFlow with matched coordinates (Section˜5.3) and evaluate it using the initialized structures described in Section˜4.2.
@@
Results. Table˜2 shows that MOFFlow-2 outperforms all baselines across both tolerance settings. Consistent with prior work [10], we observe that the optimization-based methods (RS and EA) and the general CSP baseline (DiffCSP) achieve near-zero match rates, indicating their limited effectiveness on MOF structures. Compared to MOFFlow [10], our method achieves higher match rates and lower RMSE, underscoring the importance of explicitly modeling torsion angles for accurate structure prediction. We also provide visualizations of property distributions in Section˜G.1.
Table 2: Structure prediction accuracy. We compare the structure prediction performance of random search (RS), evolutionary algorithm (EA), DiffCSP, MOFFlow, and MOFFlow-2. MR is the match rate, RMSE is the root mean squared error, and - indicates no match. $\operatorname{stol}$ is the site tolerance for matching criteria.
@@ 6.2 MOF generation
We evaluate the generative performance of MOFFlow-2 by (1) measuring validity, novelty, and uniqueness (VNU), and (2) comparing the distribution of MOF properties against that of the training set. We benchmark against MOFDiff [11], a coarse-grained diffusion model for MOF generation. Both models generate 10,000 samples using their respective pipelines. To isolate the effect of the generative model, no force-field relaxation is applied to the generated structures. Additional results on conditional generation and energy-level evaluation are provided in Section˜G.2 and Section˜G.3.
@@
Results. Table˜3 summarizes the generation performance of MOFFlow-2 compared to MOFDiff. MOFFlow-2 achieves higher scores across all metrics. Notably, MOFFlow-2 not only generates MOFs with novel combinations of known building blocks but also produces entirely new building blocks, demonstrating its potential for discovering previously unseen MOFs. In addition, MOFFlow-2 is also faster than MOFDiff, which depends on an optimization-based self-assembly procedure for final structure construction. As shown in Figure˜4, the property distribution of MOFFlow-2 aligns more closely with the ground-truth. Importantly, it spans a broader range of values than MOFDiff, indicating that MOFFlow-2 generates MOFs with more diverse physical properties. We provide visualizations of representative VNU samples generated by MOFFlow-2 in Figure˜5.
@@ 7 Conclusion
We introduced MOFFlow-2, a two-stage generative model for MOF design and structure prediction. In the first stage, the building block generator designs MOF sequences in SMILES, which are initialized into 3D structures using a predefined metal library and RDKit [14]. In the second stage, the structure prediction module assembles the complete MOF by jointly predicting rotations, translations, torsions, and lattice parameters. Experimental results show that MOFFlow-2 outperforms existing models in both generative design and structure prediction.
Despite its strong performance, MOFFlow-2 still has several limitations, including dependence on RDKit for conformer initialization, partial conditioning in property-guided generation (Section˜G.2), and approximate energy evaluation with machine learning interatomic potentials (Section˜G.3). We provide a detailed discussion of these limitations in Appendix˜H.
@@ Appendix A Implementation details for building block generator
We provide additional details for implementing the building block generator (Section˜4.1), including the tokenization process, batching strategy, and hyperparameter settings.
@@
We illustrate the tokenization process for a canonical MOF sequence defined in Section˜4.1, which transforms the string into a sequence of discrete symbols $b_{1},\dots,b_{S}]$.
@@
We then map each token to a unique vocabulary index to get $b_{1},\dots,b_{S}]$.
@@
Hyperparameters. In Table˜4, we present the hyperparameters for training the building block generator, including the batch configuration, model architecture, and optimizer settings.
Codebase. We implemented our sequence model with x-transformers [43]. We appreciate the author for the open-source implementation.
@@ Appendix B Analysis and visualization of metal building blocks
Figure˜6(a) and Figure˜6(b) show the RMSD histograms of the extracted metal building blocks for the generation and structure prediction tasks, respectively. There are only 7 and 8 distinct metal building block types, despite the large sizes of the training datasets (>150k). The RMSD histograms further indicate that all metal types exhibit low structural variability. These observations justify our template-based strategy for initializing the metal structures, described in Section˜4.2. We also provide visualizations of the metal building blocks in Figure˜7.
@@ C.1 Conditional flows and conditional vector fields
where $\exp:\mathfrak{so}(3)\rightarrow{}SO(3)$ and $\log:SO(3)\rightarrow{}\mathfrak{so}(3)$ are the exponential and logarithmic maps for $SO(3)$, respectively [45]. Since we target $\Delta q_{t\rightarrow{}1}:=q_{t}q_{1}^{\top}$ (Section˜5.1),
@@ C.3 Training and inference algorithm
Algorithm˜2 and Algorithm˜3 present the training and inference algorithms for the structure prediction model, respectively. During training, the model outputs rotations $\Delta\hat{\bm{Q}}_{t\rightarrow 1}\in\mathbb{R}^{M\times 3\times 3}$ and torsions $\hat{\bm{\Phi}}_{1}\in\mathbb{R}^{P\times 2}$ in the Euclidean space, which are directly supervised with the targets $\Delta\bm{q}_{t\rightarrow 1}$ and $\bm{\phi}_{1}$. At inference, these outputs are projected onto their respective manifolds as $\Delta\hat{\bm{q}}_{t\rightarrow 1}=\operatorname{Procrustes}(\Delta\hat{\bm{Q}}_{t\rightarrow 1})\in SO(3)^{M}$ and $\hat{\bm{\phi}}_{1}=\hat{\bm{\Phi}}_{1}/\|\hat{\bm{\Phi}}_{1}\|_{2}\in\mathbb{T}^{P}$.
@@
3:  Sample $\Delta\bm{q}_{1\rightarrow{}0}$, $\bm{\tau}_{0}$, $\bm{\phi}_{0}$, $\bm{\ell}_{0}$ from prior distributions defined in Section˜5.1.
4:  Interpolate $\Delta\bm{q}_{1\rightarrow{}t},\bm{\tau}_{t},\bm{\phi}_{t},\bm{\ell}_{t}$ according to Section˜C.1
5:  Apply $\bm{X}_{t}\leftarrow{}(\Delta\bm{q}_{1\rightarrow{}t},\bm{\tau}_{t},\bm{\phi}_{t})\cdot\bm{X}_{1}$ according to Section˜C.2.
@@
7:  Optimize loss $\mathcal{L}(\theta)$ in Equation˜5.
@@
1:Initialize structure $\tilde{\bm{X}}_{0}\leftarrow{}\operatorname{Initialize}(\mathcal{B}_{\text{2D}})$ according to Section˜4.2.
2:Sample $\Delta\bm{q}_{0}$, $\bm{\tau}_{0}$, $\bm{\phi}_{0}$, $\bm{\ell}_{0}$ from prior distributions defined in Section˜5.1.
@@ D.1 Initialization module
Building block index. We define the building block index $k$ as introduced in Section˜5.2 to resolve ambiguity between building blocks of the same type – i.e., those sharing the same SMILES representation. To assign consistent indices, we apply the following lexicographic ordering rules:
@@ D.2 Interaction module
The interaction module follows a Transformer encoder architecture [47] with root mean square layer normalization [RMSNorm; 48] (Figure˜2). Here, we provide further implementation details, including the model architecture (Algorithm˜4) and associated hyperparameters (Table˜5).
@@
[…]
@@ D.3 Output module
We describe the architectures of the four prediction heads in the output module (Section˜5.2). We begin with the block attention pooling module, which aggregates atom-level node embeddings from the interaction module into block-level embeddings (Algorithm˜5). We then detail the architectures of the rotation (Algorithm˜6), translation (Algorithm˜7), lattice (Algorithm˜8), and torsion heads (Algorithm˜9). The corresponding hyperparameters are summarized in Table˜6.
@@
1:# Construct building block coordinates $x_{\text{bb}}^{(m)}]_{m=1}^{M}\in\mathbb{R}^{M\times 3}$ and features $h_{\text{bb}}^{(m)}]_{m=1}^{M}\in\mathbb{R}^{M\times D}$.
@@
8:Construct edge distance features $e_{mj}]_{(m,j)\in\mathcal{E}_{\text{bb}}}$ where $e_{mj}=\operatorname{RBF}({\|x_{\text{bb}}^{(m)}-x_{j}\|_{2}})$.
@@
2:Average and MLP $\tilde{\bm{\ell}}_{1}\leftarrow{}\operatorname{MLP}(\sum_{m=1}^{M}h_{\text{bb}}^{(m)})$. $\triangleright$ Two-layer MLP with GELU activation.
@@
1:# Construct rotatable bond coordinates $x_{\text{rot}}^{(p)}]_{p=1}^{P}\in\mathbb{R}^{P\times 3}$ and features $h_{\text{rot}}^{(p)}]_{p=1}^{P}\in\mathbb{R}^{P\times D}$.
@@
9:Construct edge distance features $e_{pj}]_{(p,j)\in\mathcal{E}_{\text{rot}}}$ where $e_{pj}=\operatorname{RBF}(\|x_{\text{rot}}^{(p)}-x_{j}\|_{2})$.
@@
18:Apply $\bm{\Phi}_{1}\leftarrow{}\operatorname{MLP}(H_{\text{rot}})$ where $H_{\text{rot}}=[h_{\text{rot}}^{(p)}]_{p=1}^{P}.$ $\triangleright$ Two-layer MLP with GELU activation.
@@ Appendix E MOF matching
We describe the MOF matching procedure, a preprocessing step used to reduce distributional shift between training and inference, as outlined in Section˜5.3. Given an MOF structure and its building blocks, we first determine whether each block contains metal elements to classify it as metal or organic. For metal blocks, we retrieve the corresponding template from the metal library and align it to the ground-truth structure. For organic blocks, we initialize the structure using RDKit, optimize torsion angles via differential evolution to minimize RMSD with the ground-truth [16], and then align the result. After processing all blocks, we compute the RMSD between the reconstructed and ground-truth structures. This procedure is repeated three times, and structures with final RMSD below 0.5Å are retained. The full algorithm for a single matching iteration is shown in Algorithm˜10.
@@
[…]
@@
4:   Retrieve the corresponding template structure $\tilde{X}^{(m)}$ from metal library (Section˜4.2).
@@
8:   Initialize structure $\tilde{X}^{(m)}$ with RDKit (Section˜4.2).
@@ Appendix F Experimental details
Data preprocessing. We detail the data preprocessing steps described in Section˜6. First, we discard MOF structures with more than 20 building blocks since large structures that are too large may be difficult to synthesize [11]. Next, we extract key features from each structure, including Cartesian coordinates, RDKit-derived atomic features, Niggli-reduced cells, symmetrically equivalent coordinates for each building block (Section˜5.3), and canonical atoms defining the torsion angles (Section˜5.3). We then construct the metal library (Section˜4.2) and perform MOF matching (Appendix˜E) on the training dataset. Finally, since the dataset is synthetic [36], we filter out invalid MOFs using MOFChecker [38], which ensures the presence of key elements (e.g., C, H, and a metal), no atomic overlaps or coordination issues, sufficient porosity, and the absence of highly charged fragments or isolated molecules.
Data statistics. We present the data statistics for structure prediction (Section˜6.1) in Tables˜7, 8 and 9 and those for generation (Section˜6.2) in Tables˜10 and 11.
Property (number of samples =157,474=157,474) Min Mean Max
@@
Property (number of samples =19,603=19,603) Min Mean Max
@@
Property (number of samples =19,792=19,792) Min Mean Max
@@
Property (number of samples =187,047=187,047) Min Mean Max
@@
Property (number of samples =10,243=10,243) Min Mean Max